If 32 mL

mL of 5.4 M

M H

2

SO

4

H2SO4 was spilled, what is the minimum mass of NaHCO

3

NaHCO3 that must be added to the spill to neutralize the acid?


1
Expert's answer
2021-10-20T02:53:43-0400

Given Molarity of H2SO4 = 5.4 M

Volume of H2SO4 = 32 mL


the balanced chemical equation of the reaction will be

H2SO4 + 2NaHCO3 --> Na2SO4 + 2CO2 + 2H2O


Molarity = (Moles * 1000) / Volume

5.4 = ((Moles of H2SO4 ) * 1000) / 32

Moles of H2SO4 = (5.4 * 32)/1000 = 0.1728 moles


Now from balanced chemical reaction, we can see that 1 mole of H2SO4 reacts with 2 moles of NaHCO3

Hence, moles of NaHCO3 = 2 *0.1728 = 0.3456 moles


Molecular mass of NaHCO3 = 23 + 1 + 12 + (3*16) = 84 g/mole


So, mass of NaHCO3 = Moles NaHCO3 * Molecular mass

mass of NaHCO3 = 84 * 0.3456 = 29g


So the minimum amount of NaHCO3 that must be added to neutralize the acid is 29g.